5.A loan of Rs. 30,000 is to be repaid with interest at 5.41% p.a. compounded 2-monthly in equal installments payable at the beginning of every 6 months for 10 years. What is the size of each installment? What amount will be paid as total interest for this loan?

6.A person deposits Rs. 2,000 p.a. for 7 years, the first deposit is made at year zero. After 7 years the quarterly deposits of Rs. 750 are made for the next 4 years. What is the amount payable to him/her immediately after the last payment, if the rate of interest if 4% p.a.? What’s the PV for the same?

7.Find the amount of an annuity due of Rs. 1,850 payable at the for 15 years, if the interest is compounded at a rate of 7.5% effective for 1o years and 8% effective thereafter.

8.A man borrows Rs. 80,000 at compound interest of 5.68% per annum and agrees to repay the money in 15 equal 4-monthly instalments. What should be the amount of each installment? Assume that the first payment is made at the end of 5 years.



1
Expert's answer
2021-05-07T09:16:59-0400

5.

"PMT= \frac{P\times r}{1-(1+r)^{-nt}}"

Where r is rate, P is principal and t is time.

"PMT= \frac{30,000}{(1-(1+0.02705)^{-10\times0.02705})/r}=7,978.35"


6.After 7 years.

"Amount=\frac{2000(1.04)^{7}}{0.04}=68,428"


Compound this for 4 years using compound interest fo rmula

"68428\times0.04\times4+68428=79,376"


750 quarterly for 4 years

"=\frac{750\times(1.01)^{16}}{0.01}\times1.01=88,823"


Total = 88823+79376=168,199